OVHCloud Web Hosting Status

Current status
Legend
  • Operational
  • Degraded performance
  • Partial Outage
  • Major Outage
  • Under maintenance
FS#7746 — Optical loop: Paris little ring
Incident Report for Web Cloud
Resolved
We have a fiber cut on the loop of Paris small ring


The service is not interrupted.

Update(s):

Date: 2012-12-11 21:13:31 UTC
The welds seem to be done. The secure is up.

Date: 2012-12-11 21:13:04 UTC
The box is set up. The welds begin...

Date: 2012-12-11 21:12:41 UTC
A new box is being installed.

Date: 2012-12-11 11:32:01 UTC
Welds will start.

Date: 2012-12-10 22:53:52 UTC
Latest info, 2 cut-offs including one due to the Civil Engineering.

Welds haven't began yet.
Reparation won't be done before tomorrow.


Date: 2012-12-10 22:51:20 UTC
Technicians are taking some measures in order to find out all cut-off levels.

Date: 2012-12-10 14:18:25 UTC
Technicians are into place.

Date: 2012-12-10 12:34:32 UTC
The interruption has been identified.
The problem is located on a portion of SNCF track.
The persons entitled to intervene near the
rails are starting .

For the moment no ETA. However the reparation is scheduled for today.

Date: 2012-12-10 12:32:19 UTC
A general problem of fiber cut in SFR.
They are working on the problem.
Posted Dec 10, 2012 - 09:35 UTC